### v1

@@ 3.1 Asymptotic minors
Now observe that the $(b_{k},y)$-subpath of $P$ has at least $24K+$ 2 antichain vertices. Therefore, Proposition 2 implies that $\mathsf{d}\left(y,b_{k}\right)\geq 4K+1+|\mathsf{d}\left(r,b_{k}\right)-\mathsf{d}\left(r,y\right)|$. But this is is a contradiction. ∎
Claim 5.9.
@@
This follows from 5.6 and the definition of $Z_{i}$. ∎
Now we show that the branch sets constructed above are sufficiently far away from the branch paths as required.
Claim 5.10.
For any $i\in[t-1]$, $\mathsf{d}\left(D_{i},T\right)\geq K$.
@@
Suppose there is a vertex $x\in D_{i}$ and $y\in T$ such that $\mathsf{d}\left(x,y\right)<K$. The definition of $T$ implies that $y\in V(Q_{j})$ for some $j\in[t]$. Since the $(b_{j},x)$-subpath of $P$ contains at least $34K+$ 3 antichain vertices, we have a contradiction due to 5.4. ∎
Using similar arguments as above, we can prove the following.
Claim 5.11.
For any $i\in[t-1],j\in[t]$ with $(j-i)\notin\{0,1\}$, we have $\mathsf{d}\left(D_{i},Z_{j}\right)\geq K$.
Completion of the proof of the Lemma 2: Now consider the sets $T,Z_{1},\ldots,Z_{t}$, $\mathcal{Q}=\{Q^{\prime}_{1},Q^{\prime}_{2},\ldots,Q^{\prime}_{t}\}$, and $\mathcal{D}=\{D_{1},D_{2},\ldots,D_{t-1}\}$. Observe that
  • for each $i\in[t]$, $T\cap Q^{\prime}_{i}\neq\emptyset$;
  • $D_{1}\cap Z_{1}\neq\emptyset,D_{t-1}\cap Z_{t}\neq\emptyset$; and
Now 5.7, 5.8, 5.9, 5.10, and 5.11 imply that the sets $T,Z_{1},\ldots,Z_{t}$ and $\mathcal{Q}\cup\mathcal{D}$ form a $K$-fat minor model of $U_{t}$, a contradiction.
@@ 3.2 Monoholed graphs
Now, we shall prove Theorem 2. We need the definitions of Truemper configurations [34]. A theta is a graph made of three vertex-disjoint induced paths $P_{1}=a\ldots b$, $P_{2}=a\ldots b$, $P_{3}=a\ldots b$ of lengths at least 2, and such that no edges exist between the paths except the three edges incident to $a$ and the three edges incident to $b$. A pyramid is a graph made of three induced paths $P_{1}=a\ldots b_{1}$, $P_{2}=a\ldots b_{2}$, $P_{3}=a\ldots b_{3}$, two of which have length at least $2$, that are vertex-disjoint (except at $a$), such that $b_{1}b_{2}b_{3}$ is a triangle, and no edges exist between the paths except those of the triangle and the three edges incident to $a$. A prism is a graph made of three vertex-disjoint induced paths $P_{1}=a_{1}\ldots b_{1}$, $P_{2}=a_{2}\ldots b_{2}$, $P_{3}=a_{3}\ldots b_{3}$ of lengths at least $1$, such that $a_{1}a_{2}a_{3}$ and $b_{1}b_{2}b_{3}$ are triangles and no edges exist between the paths except those of the two triangles.
@@
Figure 4: (a)The graph $U_{4}$. (b) Illustration of the proof of Theorem 2. The thick, black, closed curve indicates the induced cycle $D$ (in 5.12).
Assume for contradiction that $G$ contains a $K$-fat minor model of $U_{4}$ with $K\geq 10$. Let the vertices of $U_{4}$ be denoted by $a_{0},a_{1},a_{2},a_{3},a_{4}$ as shown in Figure 4(a). Unless stated explicitly, all addition operations in this proof will be taken modulo $5$. Let $B_{i}$ denote the branch set corresponding to the vertex $a_{i}$ and $C_{i}^{j}$ denote the branch path corresponding to the edge between $a_{i}$ and $a_{j}$. See Figure 4(b) for illustrations. For the sake of readability, we shall use the same notations to indicate the subgraphs induced by the branch sets and branch paths. Let $Z=\{(0,1),(0,2),(0,3),(0,4),(1,2),(2,3),(3,4)\}$. A witness of a branch path $C_{i}^{j},(i,j)\in Z$ is an induced path $P$ such that $V(P)\subseteq C_{i}^{j}$, one end-vertex of $P$ lies in $B_{i}$, and the other end-vertex of $P$ lies in $B_{j}$. For $0\leq i\leq 4$, let $W_{i}$ be a witness of $C_{i}^{i+1}$ of shortest possible length, and let $u_{i}\in B_{i},v_{i+1}\in B_{i+1}$ be the end-vertices of $W_{i}$. For $0\leq i\leq 4$, $X_{i}$ denote a path of shortest possible length between $u_{i}$ and $v_{i}$ such that $V(X_{i})\subseteq B_{i}$. Let $H$ be the subgraph of $G$ induced by $\bigcup\limits_{i=0}^{4}V(X_{i})\cup V(W_{i})$.
Claim 5.12.
The graph $H$ contains an induced cycle $D$ such that for each $0\leq i\leq 4$, both $V(B_{i})\cap V(D)$ and $V(C_{i}^{i+1})\cap V(D)$ induce paths of lengths at least $K-2$.
@@
The minimality of $W_{i}$ implies that only the end-vertices of $W_{i}$ or their neighbors (in $W_{i}$) can be adjacent to some vertices of $X_{i}$ or $X_{i+1}$. Hence, there must exist a subpath of $W_{i}$, say $W_{i}^{\prime}$, of length at least $K-2$ such that no internal vertex of $W_{i}^{\prime}$ is adjacent to a vertex of $X_{i}$ or $X_{i+1}$. Since the distance between two witnesses (corresponding to two branch paths) is at least $K$, observe that there must exist a subpath of $X_{i}$, say $X^{\prime}_{i}$, of length at least $k-2$ none of whose internal vertex is adjacent to the vertices of $W_{i}$ or $W_{i+1}$. Now, using $W_{i}^{\prime}$ and $X^{\prime}_{i}$, $0\leq i\leq 4$, it is possible to construct an induced cycle $D$ satisfying the claim. ∎
Let $D$ be the induced cycle whose existence was proved in 5.12. For $i\in\{0,1,2,3,4\}$, let $D_{i}=B_{i}\cap D$. Let $Q$ be an induced path of the shortest possible length between a vertex of $D_{2}$ and $D_{0}$ such that $V(Q)\subseteq B_{0}\cup C_{0}^{2}\cup B_{2}$. (Since $B_{0},C_{0}^{2},B_{2}$ are connected induced subgraphs, $Q$ always exists). Observe that $Q$ and $D$ together creates a balanced $3$-path configuration $P_{1}$, $P_{2}$, $P_{3}$ such that $V(P_{1})\cap V(W_{0})\neq\emptyset,V(P_{2})\cap C_{0}^{2}\neq\emptyset$ and $V(P_{3})\cap V(W_{4})\neq\emptyset$. For $i\in[3]$, let $z_{i}$ (resp. $z^{\prime}_{i})$ denote the end-vertex of $P_{i}$ which is adjacent to some vertices of $D_{0}$ (resp. $D_{2})$. Let $K^{\prime}$ be the lengths of the three paths. Observe that the vertices of the paths $P_{2}$ and $P_{1}$ (resp. $P_{3}$) induces a cycle $F$ (resp. $F^{\prime}$) of order at most $22K^{\prime}+$ 2 which also contains $D_{1}$ (resp. $D_{3}$).
Now consider an induced path $R$ of the shortest possible length between a vertex of $D_{3}$ and $D_{0}$ such that $V(R)\subseteq B_{0}\cup C_{0}^{3}\cup B_{3}$. Let $x$ and $y$ denote the end-vertices of $R$ that lie on $D_{3}$ and $D_{0}$, respectively. Let $x^{\prime},y^{\prime}$ be the neighbours of $x$ and $y$, respectively and $R^{\prime}$ be the $(x^{\prime},y^{\prime})$-subpath of $R$. The minimality of $R$ implies that no internal vertex of $R^{\prime}$ is adjacent to any vertex of $D_{0}$ and $D_{3}$. Furthermore, we prove the following claim.
Claim 5.13.
@@
Suppose $w$ is the vertex of $R^{\prime}$, which is closest to $x^{\prime}$ and is also adjacent to some vertex $w^{\prime}$ of $F^{\prime}$. If $w^{\prime}\in P_{3}$, then $F^{\prime}$ together with the $(w,x^{\prime})$-subpath must create a $3$-path configuration which is either unbalanced or has length strictly less than $K^{\prime\prime}$, a contradiction. For similar reasons, $w^{\prime}$ cannot be $z_{2}$. Now consider the case when $w^{\prime}$ is an internal vertex of $P_{2}$. Let $U$ be the maximal subpath of $P_{3}$ whose one end-vertex is $z^{\prime}_{3}$, and no internal vertex of $U$ is a neighbor of $x^{\prime}$. Observe that $U$ together with the $(w,x^{\prime})$-subpath of $R^{\prime}$ creates an induced path $U^{\prime}$ whose (only) end-vertices are adjacent to the vertices of $F$ (the induced cycle created by $P_{2}$ and $P_{1}$). Since $w^{\prime}$ is an internal vertex, $F$ together with $U^{\prime}$ must create a $3$-path configuration which is either unbalanced or has a length strictly less than $K^{\prime}$, a contradiction. ∎
Let $S$ be the $(y,z_{3})$-subpath of $D_{0}$ and let $S^{\prime}$ be the subpath of $S$ of minimum possible length such that one end-vertex of $S^{\prime}$ is adjacent to $y^{\prime}$ and the other end-vertex is adjacent to $z_{2}$. Let $R^{\prime\prime}$ denote the path induced by $V(R^{\prime})\cup V(S^{\prime})$. The induced cycle $F^{\prime}$ along with $R^{\prime\prime}$ creates an unbalanced $3$-path configuration, a contradiction. ∎

### v2

@@ 3.1 Asymptotic minors
If $z=b^{\prime}_{k}$, then let $z^{\prime}$ be the vertex on $Q_{i}$ such that $\mathsf{d}\left(r,b^{\prime}_{k}\right)=\mathsf{d}\left(r,z^{\prime}\right)$. From $\mathsf{d}\left(d_{i},z\right)=\mathsf{d}\left(d^{\prime}_{i},b^{\prime}_{k}\right)\leq K$, $\mathsf{d}\left(r,b^{\prime}_{k}\right)\leq\mathsf{d}\left(r,d^{\prime}_{i}\right)$, and $\mathsf{d}\left(d^{\prime}_{k},d^{\prime}_{i}\right)\leq K$, it follows that $\mathsf{d}\left(z^{\prime},d^{\prime}_{i}\right)\leq K$. Hence,
$\mathsf{d}\left(b_{k},y\right)$ $\leq 1+\mathsf{d}\left(z,d^{\prime}_{k}\right)+\mathsf{d}\left(d^{\prime}_{k},d^{\prime}_{i}\right)+\mathsf{d}\left(z^{\prime},d^{\prime}_{i}\right)+\mathsf{d}\left(z^{\prime},y^{\prime}\right)+\mathsf{d}\left(y^{\prime},x\right)+\mathsf{d}\left(x,y\right)$
$\leq 1+(\mathsf{d}\left(r,z^{\prime}\right)-\mathsf{d}\left(r,y^{\prime}\right))+\mathsf{d}\left(b^{\prime}_{k},d^{\prime}_{k}\right)+\mathsf{d}\left(d^{\prime}_{k},d^{\prime}_{i}\right)+\mathsf{d}\left(y^{\prime},x\right)+\mathsf{d}\left(x,y\right)+\mathsf{d}\left(z^{\prime},y^{\prime}\right)$
$<1+\mathsf{d}\left(r,b^{\prime}_{k}\right)-\mathsf{d}\left(r,y\right)+5K$
The above arguments imply that $\mathsf{d}\left(b_{k},y\right)\leq 5K+|\mathsf{d}\left(r,b_{k}\right)-\mathsf{d}\left(r,y\right)|$. Now, observe that the $(b_{k},y)$-subpath (say $P^{\prime}$) of $P$ contains $D_{k}$. Therefore, $P^{\prime}$ contains at least one component of the graph obtained by removing the edges of $D_{k}$ fromn the $(b_{k},b_{k+1})$-subpath of $P$. The definition of $D_{k}$ now implies that the $(b_{k},y)$-subpath of $P$ has at least $5K+2$ antichain vertices of $A$. Therefore, Proposition˜13 implies that $\mathsf{d}\left(y,b_{k}\right)\geq 5K+1+|\mathsf{d}\left(r,b_{k}\right)-\mathsf{d}\left(r,y\right)|$, a contradiction.
Now, consider the case where $\mathsf{d}\left(r,y\right)>\mathsf{d}\left(r,d^{\prime}_{i}\right)$: see Figure˜7(b). Let $y^{\prime}$ be the vertex of $Q_{k}$ such that $\mathsf{d}\left(r,y^{\prime}\right)=\min\{\mathsf{d}\left(r,b^{\prime}_{k}\right),\mathsf{d}\left(r,y\right)\}$. Let $x^{\prime}$ be the vertex of $Q_{k}$ such that $\mathsf{d}\left(r,x\right)=\mathsf{d}\left(r,x^{\prime}\right)$. Since $\mathsf{d}\left(x,y\right)<K$, $\mathsf{d}\left(r,x\right)=\mathsf{d}\left(r,x^{\prime}\right)$, and $x^{\prime}$ lies on a $(r,y^{\prime})$-isometric path, we have $\mathsf{d}\left(x^{\prime},y^{\prime}\right)<K$. Since $\mathsf{d}\left(r,x\right)<\mathsf{d}\left(r,d^{\prime}_{i}\right)<\mathsf{d}\left(r,y\right)$ and $\mathsf{d}\left(x,y\right)<K$, we have $\mathsf{d}\left(x,d^{\prime}_{i}\right)<K$. Now observe the following:
$\mathsf{d}\left(b_{k},y\right)$ $\leq 1+\mathsf{d}\left(b^{\prime}_{k},y^{\prime}\right)+\mathsf{d}\left(x^{\prime},y^{\prime}\right)+\mathsf{d}\left(d^{\prime}_{k},x^{\prime}\right)+\mathsf{d}\left(d^{\prime}_{k},d^{\prime}_{i}\right)+\mathsf{d}\left(d^{\prime}_{i},x\right)+\mathsf{d}\left(x,y\right)$
$\leq 1+|\mathsf{d}\left(r,b^{\prime}_{k}\right)-\mathsf{d}\left(r,y\right)|+\mathsf{d}\left(x^{\prime},y^{\prime}\right)+\mathsf{d}\left(d^{\prime}_{k},x^{\prime}\right)+\mathsf{d}\left(d^{\prime}_{k},d^{\prime}_{i}\right)+\mathsf{d}\left(d^{\prime}_{i},x\right)+\mathsf{d}\left(x,y\right)$
$\leq 1+|\mathsf{d}\left(r,b^{\prime}_{k}\right)-\mathsf{d}\left(r,y\right)|+5K$
Now, observe again that the $(b_{k},y)$-subpath of $P$ has at least $5K+2$ antichain vertices of $A$. Therefore, Proposition˜13 implies that $\mathsf{d}\left(y,b_{k}\right)\geq 5K+1+|\mathsf{d}\left(r,b_{k}\right)-\mathsf{d}\left(r,y\right)|$. But this is a contradiction.
Now we prove Item˜7.
Claim 3.1.I.
@@
The definitions of $D_{i}$ and $D_{j}$ imply that both of them contain at least 2K22K-2 antichain vertices each, implying the length of both $D_{i}$ and $D_{j}$ are at least 2K32K-3. Since $K\geq 5$, we have 2K3>K+12K-3>K+1. Now, the claim follows from the definitions of $Z_{i}$ and $Z_{j}$. ∎
Now we show that the constructed branch sets are sufficiently far away from the branch paths, as required. First, we prove Item˜8.
Claim 3.1.J.
For any $i\in[t-1]$, $\mathsf{d}\left(T,D_{i}\right)\geq K$.
@@
Suppose that there is a vertex $x\in D_{i}$ and $y\in T$ such that $\mathsf{d}\left(x,y\right)<K$. The definition of $T$ implies that $y\in V(Q_{j})$ for some $j\in[t]$. Since the $(b_{j},x)$-subpath of $P$ contains at least $5K+2$ antichain vertices of $A$, we have a contradiction due to ˜3.1.E. ∎
Using similar arguments as ˜3.1.J, we can prove Item˜9.
Claim 3.1.K.
For any $i\in[t-1],j\in[t]$ with $|j-i|\notin\{0,1\}$, we have $\mathsf{d}\left(D_{i},Z_{j}\right)\geq K$.
The proof of Theorem˜14 is now complete.
@@ 3.2 Monoholed graphs
Now, we shall prove Theorem˜6. We need the definitions of Truemper configurations [47]. A theta is a graph made of three internally vertex-disjoint induced paths $P_{1}=a\ldots b$, $P_{2}=a\ldots b$, $P_{3}=a\ldots b$ of lengths at least 2, and such that no edges exist between the paths except the three edges incident to $a$ and the three edges incident to $b$. A pyramid is a graph made of three induced paths $P_{1}=a\ldots b_{1}$, $P_{2}=a\ldots b_{2}$, $P_{3}=a\ldots b_{3}$, two of which have length at least $2$, that are vertex-disjoint (except at $a$), such that $b_{1}b_{2}b_{3}$ is a triangle, and no edges exist between the paths except those of the triangle and the three edges incident to $a$. A prism is a graph made of three vertex-disjoint induced paths $P_{1}=a_{1}\ldots b_{1}$, $P_{2}=a_{2}\ldots b_{2}$, $P_{3}=a_{3}\ldots b_{3}$ of lengths at least $1$, such that $a_{1}a_{2}a_{3}$ and $b_{1}b_{2}b_{3}$ are triangles and no edges exist between the paths except those of the two triangles.
@@
Figure 8: (a) The graph $U_{4}$. (b) Illustration of the proof of Theorem˜6. The thick, black, closed curve indicates the induced cycle $D$ (in ˜3.2.L).
Assume for contradiction that $G$ contains a $K$-fat minor model of $U_{4}$ with $K\geq 10$. Let the vertices of $U_{4}$ be denoted by $a_{0},a_{1},a_{2},a_{3},a_{4}$ as shown in Figure˜8(a). Unless stated explicitly, all addition operations in this proof will be taken modulo $5$. Let $B_{i}$ denote the branch set corresponding to the vertex $a_{i}$ and $C_{i}^{j}$ denote the branch path corresponding to the edge between $a_{i}$ and $a_{j}$. See Figure˜8(b) for illustrations. For the sake of readability, we shall use the same notations to indicate the subgraphs induced by the branch sets and branch paths. Let $Z=\{(0,1),(0,2),(0,3),(0,4),(1,2),(2,3),(3,4)\}$. A witness of a branch path $C_{i}^{j},(i,j)\in Z$ is an induced path $P$ such that $V(P)\subseteq C_{i}^{j}$, one end-vertex of $P$ lies in $B_{i}$, and the other end-vertex of $P$ lies in $B_{j}$. For $0\leq i\leq 4$, let $W_{i}$ be a witness of $C_{i}^{i+1}$ of shortest possible length, and let $u_{i}\in B_{i},v_{i+1}\in B_{i+1}$ be the end-vertices of $W_{i}$. For $0\leq i\leq 4$, let $X_{i}$ denote a path of shortest possible length between $u_{i}$ and $v_{i}$ such that $V(X_{i})\subseteq B_{i}$.
Let $H$ be the subgraph of $G$ induced by $\bigcup\limits_{i=0}^{4}V(X_{i})\cup V(W_{i})$.
Claim 3.2.L.
The graph $H$ contains an induced cycle $D$ such that for each $0\leq i\leq 4$, both $V(B_{i})\cap V(D)$ and $V(C_{i}^{i+1})\cap V(D)$ induce paths of lengths at least K2K-2.
@@
The minimality of $W_{i}$ implies that only the end-vertices of $W_{i}$ or their neighbors (in $W_{i}$) can be adjacent to some vertices of $X_{i}$ or $X_{i+1}$. Recall that for each $0\leq i\leq 4$, $W_{i}$ has length at least $K$ since it connects two branch sets. Hence, there must exist a subpath of $W_{i}$, say $W_{i}^{\prime}$, of length at least K2K-2, such that no internal vertex of $W_{i}^{\prime}$ is adjacent to a vertex of $X_{i}$ or $X_{i+1}$. Similarly, since the distance between two witnesses (corresponding to two branch paths) is at least $K$, observe that there must exist a subpath of $X_{i}$, say $X^{\prime}_{i}$, of length at least K2K-2, none of whose internal vertex is adjacent to the vertices of $W_{i-1}$ or $W_{i}$. Now, using $W_{i}^{\prime}$ and $X^{\prime}_{i}$, $0\leq i\leq 4$, it is possible to construct an induced cycle $D$ satisfying the claim. ∎
Let $D$ be the induced cycle whose existence was proved in ˜3.2.L. For $i\in\{0,1,2,3,4\}$, let $D_{i}=B_{i}\cap D$. Let $Q$ be an induced path of the shortest possible length between a vertex of $D_{2}$ and $D_{0}$ such that $V(Q)\subseteq B_{0}\cup C_{0}^{2}\cup B_{2}$. Since $B_{0},C_{0}^{2},B_{2}$ are connected induced subgraphs, $Q$ always exists. Notice that $Q$ and $D$ together create a $3$-path configuration (which is balanced since $G$ is monoholed), say $P_{1}$, $P_{2}$, $P_{3}$, such that $V(P_{1})\cap V(W_{0})\neq\emptyset,V(P_{2})\cap C_{0}^{2}\neq\emptyset$ and $V(P_{3})\cap V(W_{4})\neq\emptyset$. See Figure˜8(b) for an illustration. For $i\in[3]$, let $z_{i}$ (resp. $z^{\prime}_{i})$ denote the end-vertex of $P_{i}$ which is adjacent to some vertices of $D_{0}$ (resp. $D_{2})$. Let $K^{\prime}$ be the length of the three paths. Observe that the vertices of the paths $P_{2}$ and $P_{1}$ (resp. $P_{2}$ and $P_{3}$) are contained in a cycle $F$ (resp. $F^{\prime}$) of order at most $2K^{\prime}+2$ which also contains $D_{1}$ (resp. $D_{3}$). ote that the vertices of $F$ are either contained in $P_{1}$ or $P_{2}$. Similarly, the vertices of $F^{\prime}$ are contained in either $P_{2}$ or $P_{3}$. Moreover, $D_{3}$ is a subpath of $P_{3}$.
Now, consider an induced path $R$ of the shortest possible length between a vertex of $D_{3}$ and $D_{0}$ such that $V(R)\subseteq B_{0}\cup C_{0}^{3}\cup B_{3}$. Let $x$ and $y$ denote the end-vertices of $R$ that lie on $D_{3}$ and $D_{0}$, respectively. Let $x^{\prime},y^{\prime}$ be the neighbours of $x$ and $y$ on $R$, respectively and $R^{\prime}$ be the $(x^{\prime},y^{\prime})$-subpath of $R$. The minimality of $R$ implies that no internal vertex of $R^{\prime}$ is adjacent to any vertex of $D_{0}$ and $D_{3}$. Furthermore, we prove the following claim.
Claim 3.2.M.
@@
Suppose for a contradiction that there is a vertex of $R^{\prime}$ which is adjacent to some vertex of $F^{\prime}$. Let $w$ be the vertex of $R^{\prime}$, which is closest to $x^{\prime}$ and is also adjacent to some vertex $w^{\prime}$ of $F^{\prime}$. Note that $w^{\prime}\in V(P_{2})$ or $w^{\prime}\in V(P_{3})$.
Suppose first that $w^{\prime}\in V(P_{3})$. Observe that $x$ is a vertex of $D_{3}$, which implies that $x\in V(P_{3})$. Let $P^{\prime}$ denote the path created by the vertices of the $(w,x^{\prime})$-subpath of $R$ along with $w^{\prime}$ and $x$. Observe that both endpoints of $P^{\prime}$ lie in $P_{3}$. Therefore, the $(w^{\prime},x)$-subpath of $P_{3}$ has strictly smaller length than $P_{3}$. Hence, $F^{\prime}$ together with $P^{\prime}$ must create a $3$-path configuration, which is either unbalanced or has length strictly less than $K^{\prime}$, a contradiction. For similar reasons, $w^{\prime}$ cannot be $z_{2}$.
Now, consider the case where $w^{\prime}$ is an internal vertex of $P_{2}$. Let $U$ be the maximal subpath of $P_{3}$ whose one end-vertex is $z^{\prime}_{3}$, and no internal vertex of $U$ is a neighbor of $x^{\prime}$. Observe that $U$ together with the $(w,x^{\prime})$-subpath of $R^{\prime}$ creates an induced path $U^{\prime}$ whose (only) end-vertices are adjacent to the vertices of $F$ (the induced cycle created by $P_{2}$ and $P_{1}$). In particular, both end-vertices of $U^{\prime}$ lies on $P_{2}$ or is adjacent to a vertex of $P_{2}$. Since $w^{\prime}$ is an internal vertex of $P_{2}$, $F$ together with $U^{\prime}$ must create a $3$-path configuration $Q$ containing a path which is a proper subpath of $P_{2}$. Hence, either $Q$ is unbalanced or has a length strictly less than $K^{\prime}$, a contradiction. ∎
Let $S$ be the $(y,z_{3})$-subpath of $D_{0}$ and let $S^{\prime}$ be the subpath of $S$ of minimum possible length such that one end-vertex of $S^{\prime}$ is adjacent to $y^{\prime}$ and the other end-vertex is adjacent to $z_{2}$. Let $R^{\prime\prime}$ denote the path induced by $V(R^{\prime})\cup V(S^{\prime})$. Then, $F^{\prime}$ along with $R^{\prime\prime}$ contains a $3$-path configuration $Q$ which is either unbalanced or has length strictly greater than $K^{\prime}$, a contradiction. ∎